# High-throughput Parasitic-independent Probe Thermal Resistance Calibration for Robust Thermal Mapping with Scanning Thermal Microscopy

### v2

@@ III.4 Modeling contact thermal resistance ()
The interfacial thermal resistance () is determined using FEM modeling, as discussed in Section IV. Furthermore, heat-spreading analysis is crucial for determining thermal conductance through the sample. Yovanovich et al.yovanovich2002analytical and Muzychka et al.muzychka1999thermal derived analytical solutions for the thermal spreading resistance (), as shown in Eq. (13), for a film on a substrate as a function of the film thickness () and the thermal conductivities of the film () and substrate ().
where is defined as , is the source radius, and is the first-order Bessel function of the first kind. Furthermore, Muzychka and Yovanovich muzychka2004thermal derived a more compact expression for the spreading resistance of a circular isoflux heat source on a finite-thickness isotropic slab. Instead of evaluating the full integral in Eq. 13, the spreading resistance can be expressed in a closed correlation form as
@@ IV.1 Finite-element multiphysics modeling of SThM probe
The variation in the working environment of SThM measurements and the uncertainty in the thermal resistance of the tip-sample interface () make quantitative measurements challenging. Therefore, a fundamental understanding of thermal transport at the tip-sample interface is essential for the quantitative determination of local thermal conductivities of the sample. In this study, a detailed three-dimensional finite element model (3D-FEM) is developed to evaluate the interfacial thermal resistance () by simulating the heat transfer mechanisms occurring at the probe–sample junction. Accurate determination of provides insights into the complex heat transfer pathways occurring between the tip and the sample. This 3D-FEM model is developed using COMSOL Multiphysics® to systematically explore the effects of multiple physical parameters on the tip–sample thermal interaction. Fig. 5(a) illustrates the SThM probe and sample structure, along with their meshed geometry. The probe comprises a cantilever, an pad connected to the tip, and a tip. Detailed dimensions and structural specifications of probe are provided in Fig. S5 (Supplementary Material). The model incorporates two primary physical processes: (1) heat transfer in solids and (2) electromagnetic heating (Joule heating in the resistor). All exposed surfaces are set as adiabatic, except for the interface between the tip and the sample.
@@
The quasi-equilibrium heat transfer equation with a Joule heat source is used to solve the heat transfer throughout the system:
Here, , , and denote the thermal conductivity, density, and heat capacity at constant pressure, respectively, for each domain of the FEM model. is the Joule heat generated in the SThM tip when is applied across the probe.
@@
The heat generated at the probe tip is transferred through the tip–sample interface and subsequently into the sample. To accurately capture the variation in , the SThM tip–sample contact is modeled over a wide range of values, from to K/W, as illustrated in Fig. 4(a). This simulation range is selected based on experimentally obtained thermal resistance values for nanoscale contacts. At low interfacial resistance ( K/W), the contribution of to the overall thermal resistance () is negligible, indicating efficient heat transfer across the interface. In contrast, at high interfacial resistance ( K/W), heat transfer between the tip and sample becomes significantly suppressed, effectively isolating the thermal transport pathways. The resulting probe temperature profiles for different values are presented in Fig. 4(a). At high interface thermal resistance values (), the sample temperature remains constant at room temperature, indicating the absence of local heating. Therefore, cannot exceed , since local heating of the sample is observed during the measurement. Furthermore, The difference between the SThM tip temperature and the corresponding local sample temperature variation is plotted in Fig. 4(b). The change in tip and sample temperature () is shown for various values. These results demonstrate the significant influence of on the quantitative determination of the sample’s local thermal conductivity.
The effect of on local sample heating was examined for Cu (thermal conductivity: 300 Wm-1K-1) and SiO2 (thermal conductivity: 1.4 Wm-1K-1). The results shown in Fig. 5(a–h) indicate that is largely independent of the material type, consistent with the experimental observations discussed in Section IV(B). In practice, the interfacial roughness between the probe and the sample can influence the effective value; however, in the present simulations, a perfectly smooth interface is assumed. Therefore, the influence of surface roughness on is considered negligible. At lower interfacial thermal resistance ( K/W), heat transfer from the probe tip is efficiently distributed across the surrounding region of the sample surface, resulting in a broader thermal diffusion profile. In contrast, at higher interfacial resistance ( K/W), heat transfer into the sample is strongly suppressed, leading to minimal or no local heating beneath the tip. These results suggest that values exceeding K/W prevent local sample heating. In contrast, the present experimental configuration exhibits local heating, indicating that is below this threshold.
@@ IV.2 Experimental determination of model parameters
To experimentally validate the quantitative SThM model, a 15 nm aluminum (Al) thin film deposited on a SiO2 substrate was selected as the test sample. The film was patterned adjacent to the bare SiO2 substrate, enabling direct comparison between regions of differing thermal conductivity under identical measurement conditions. The sample preparation and fabrication details are provided in the corresponding section. All SThM measurements were carried out at a controlled room temperature of 294 K, monitored continuously by a temperature regulation system. The SThM probe was positioned near the Al/SiO2 interface and brought into mechanical contact with the sample surface. This contact led to a sudden reduction in probe temperature, indicating the activation of an additional thermal conduction pathway at the tip–sample junction, thereby enhancing heat dissipation into the sample (see Fig. S1 from the Supplementary Material). Subsequently, as the tip scanned the sample area, the probe recorded thermal signals from the sample as temperature variations, which allowed the calculation of the sample’s thermal resistance. A total of 7 independent SThM scans were acquired to ensure the reproducibility of the measurements. The corresponding pixel-level thermal signal and topography data were extracted using the SPM image analysis software named Gwyddion. Subsequently, the data were processed and analyzed using a custom-developed Python script designed specifically for quantitative pixel-wise thermal signal correlation and noise reduction. This in-house code enabled automated data extraction, normalization, and statistical comparison across multiple scans, significantly improving the accuracy and consistency of the thermal mapping results.
@@
Figure. 6(a,c) shows the topography and thermal signal map of 15 nm-thick Al thin film patterned on a SiO2 substrate. A strong thermal signal contrast was observed between the Al film and the SiO2 substrate, reflecting the significant differences in their thermal conductivities. Two approach-retraction curves for either side of the interface are used for the conversion between the SThM signal and thermal resistance using Equation (10). Further details about the approach-retraction curve are provided in Fig. S1 (Supplementary Material). The relative change in the thermal signal (20 mV) from the SiO2 surface to the Al film corresponds to a decrease in thermal resistance from the SiO2 surface to the 15 nm-thick Al film. Therefore, this decrease in thermal resistance indicates a higher thermal conductivity of the Al film compared to the SiO2 substrate. We assumed a literature value for the thermal conductivity of 1.38 Wm-1K-1 for the SiO2 substrate grove1967physics. Since the measured thermal resistance values are below K/W, it can be concluded that is same order of magnitude as K/W. In practice, the thermal contact between the SThM probe and the sample may slightly differ for Al and SiO2 surfaces; however, it is very challenging to measure the exact for each case. Therefore, for simplification, we assumed the same for both Al and SiO2 interfaces. This assumption introduces negligible uncertainty in the measurement because both surfaces exhibit similar roughness (see Fig. 6(a)).
@@
The thermal conductivity of Al is determined using the procedure described in Refs. huang2024violation and spiece2021quantifying, which reduces noise and measurement uncertainty. To extract the thermal resistance associated with Al and SiO2, a pixel-to-pixel correlation between the height and thermal signal maps is performed. This correlation is used to construct a 2D histogram (see Fig. S2 from the Supplementary Material), which enables a separation of the thermal responses from Al and SiO2 surfaces. Traditional analysis methods often rely on arbitrarily selected line-cuts or cropped regions of interest, introducing potential user bias in the selection process. In contrast, the present approach systematically processes all available data points from the thermal signals measurement, ensuring reproducibility and minimizing subjective interpretation. This comprehensive analysis leverages the full spatial information content, thereby improving statistical confidence in the extracted thermal properties. Subsequently, the acquired data are subjected to statistical processing to extract thermal properties. However, this analysis is susceptible to errors arising from experimental artifacts—such as surface contamination, tip-sample contact variations, or instrumental noise—which can distort the thermal signal and compromise measurement accuracy. Therefore, to quantitatively extract representative thermal resistance values while accounting for spatial heterogeneity, 2D histogram of film thickness versus thermal resistance is projected onto the resistance axis. This projection consolidates measurements across all thickness values, yielding a one-dimensional distribution from which statistically robust resistance values can be determined. The resulting 1D distribution is modeled using a double-Gaussian function, from which the mean peak positions corresponding to the thermal resistances of Al and SiO2 are extracted. The statistical uncertainties in these mean values are obtained from the covariance matrix of the nonlinear least-squares fitting procedure. From this analysis, the mean thermal resistances were found to be 2.60 0.03 MKW-1 and 5.41 0.04 MKW-1 for Al and SiO2, respectively.
@@
In the following analysis, the thermal resistance distributions presented in Figure 6 (e,f) are utilized to quantitatively determine the thermal conductivity of 15 nm Al thin film. Following the theoretical framework established in Section III, we model the measured thermal resistance as a series combination of the tip thermal resistance (), sample spreading resistance (), and the film-substrate interfacial thermal resistance (). This formulation enables direct extraction of the Al thermal conductivity from the experimental data. Therefore, the measured thermal resistance () consists of three components: , , and . Among these, ( K/W) is at least two orders of magnitude smaller than and . Moreover, since the thermal conductivity is evaluated relative to the SiO2 surface, the contribution from can be neglected. Furthermore, as established in the calibration procedure described earlier, the tip-sample contact resistance () remains constant for a given probe and is independent of the sample. Consequently, when the SThM probe contacts both the reference SiO2 surface and the Al film under identical experimental conditions, can be treated as invariant. This assumption allows us to isolate the spreading resistance component: the measured thermal resistance () on the Al surface. Finally, the thermal conductivity is calculated using Eq. 14, under the assumption of a semi-infinite substrate. The analysis of the statistical distribution of thermal resistance values yields W/m·K for the 15 nm Al film. The reported uncertainty accounts for both statistical variations and thermal signal errors arising from additional heat transfer through the water meniscus formed around the SThM tip assy2015temperature.
Discussion
@@
The thermal conductivity of Al consists of both phonon () and electron () contributions, with electrons typically dominating heat transport in bulk. Jain et al. performed detailed first-principles calculations to examine both phonon-phonon and electron-phonon scattering processes in Al, quantitatively determining the individual contributions of and to the overall thermal conductivity jain2016thermal. This study further quantified the mean free path () distributions of both phonon and electron heat carriers by analyzing their respective thermal conductivity accumulation functions at 300 K. Analysis of the thermal conductivity accumulation function demonstrates that the electron contribution decreases substantially as decreases, indicating that electrons with short contribute less to heat transport than those with longer . This suggests that long-range electron transport dominates thermal conductivity in bulk Al, whereas phonon transport shows the inverse behavior, being dominated by the short scattering processes.
@@
The accumulation function data at 300 K indicate that phonon thermal conductivity in Al is dominated by carriers with mean free paths ranging from 1 nm to 22 nm (comprising of ), while electron thermal conductivity requires longer mean free paths of 10 nm to 25 nm for similar accumulation ( of ). The onset of contributions also differs markedly: phonon transport begins at , whereas electron transport only becomes substantial at . These characteristic length scales directly impact thermal transport in confined geometries. In our 15 nm Al film, interfacial Al2O3 layer (1–2 nm thick) forms naturally due to exposure to ambient air, reducing the effective conductive metallic thickness to approximately 11–13 nm. This confinement preferentially suppresses electron transport, as the film dimension is smaller than the typical electron mean free paths, resulting in diffuse boundary scattering that significantly reduces . Conversely, phonons with mean free paths of 1–13 nm can propagate through the film with less boundary interference. This size-dependent scattering leads to a crossover in the dominant heat carrier: while bulk Al thermal conductivity is electron-dominated, phonon transport becomes the primary mechanism in this nanoscale thin film. Therefore, the thermal conductivity obtained in this study ( WmK) falls within a reasonable and expected range for nanoscale Al films. This value, approximately 80 times lower than bulk Al, is consistent with the dominance of phonon transport and the suppression of electron contributions due to dimensional confinement.
@@
The thermal signal and topography appear elevated along the Al/SiO2 boundary (see Fig. 6(a, c)). The elevated topography introduces a minor artifact in the thermal signal due to AFM feedback sensitivity effects at the step height. Consequently, the higher thermal resistance extracted at the edge arises from the same artifact, as it is directly derived from the thermal signal. Nevertheless, Fig. 6(e) demonstrates the significant potential of this approach for probing thermal transport variations across dislocation defects and grain boundaries with sub-100 nm spatial resolution.
V Conclusion

### v3

@@ III.4 Modeling contact thermal resistance ()
The impact of the interfacial thermal resistance () and its physically reasonable range are determined using FEM modeling, as discussed in Sec. IV. Furthermore, heat-spreading analysis is crucial for determining the thermal conductivity () of the sample. Yovanovich et al.yovanovich2002analytical and Muzychka et al.muzychka1999thermal derived analytical solutions for the thermal spreading resistance (), as shown in Eq. (14), for a film on a substrate as a function of the film thickness () and the thermal conductivities of the film () and substrate ().
where is defined as , is the source radius, and is the first-order Bessel function of the first kind. Furthermore, Muzychka and Yovanovich muzychka2004thermal derived a more compact expression for the spreading resistance of a circular isoflux heat source on a finite-thickness isotropic slab. Instead of evaluating the full integral in Eq. (14), the spreading resistance can be expressed in a closed correlation form as shown in Eq. (15).
@@ IV.1 Finite-element multiphysics modeling of SThM probe
The variation in the working environment of SThM measurements and the uncertainty in the thermal resistance of the tip-sample interface () make quantitative measurements challenging. Therefore, a fundamental understanding of thermal transport at the tip-sample interface is essential for the quantitative determination of local thermal conductivities of the sample. In this study, a detailed three-dimensional finite element model (3D-FEM) is developed to evaluate the interfacial thermal resistance () by simulating the heat transfer mechanisms occurring at the probe–sample junction. Accurate determination of provides insights into the complex heat transfer pathways occurring between the tip and the sample. This 3D-FEM model is developed using COMSOL Multiphysics® to systematically explore the effects of multiple physical parameters on the tip–sample thermal interaction. Fig. 4(a) illustrates the SThM probe and sample structure, along with their meshed geometry. The probe comprises a Si3N4 cantilever, an Au pad connected to the tip, and a Pd tip. Detailed dimensions and structural specifications of probe are provided in Fig. S5 (Supplementary Material). The model incorporates two primary physical processes: (1) heat transfer in solids and (2) electromagnetic heating (Joule heating in the Pd resistor). All exposed surfaces are set as adiabatic, except for the interface between the tip and the sample.
@@
The quasi-equilibrium heat transfer equation with internal Joule heating is employed to model heat transport throughout the system:
where is the thermal conductivity. The volumetric Joule heat source arises from electrical biasing of the Pd thermistor and is given by
Where is the applied voltage, is the temperature-dependent electrical resistance of the Pd thermistor, and is the effective heater volume. This expression corresponds to uniform distribution of electrical power dissipation within the resistive element. Convection term was not considered in Eqs. (16) due to its negligible contribution under the given experimental conditions zhang2023realizing; bodzenta2022scanning; deshmukh2022direct.
@@
The heat generated at the probe tip is transferred through the tip–sample interface and subsequently into the sample. To accurately capture the variation in , the SThM tip–sample contact is modeled over a wide range of values, from to K/W, as illustrated in Fig. 5(a). This simulation range is selected based on experimentally obtained thermal resistance values for nanoscale contactsli2025quantitative; gomes2015scanning. At low interfacial resistance ( K/W), the contribution of to the overall thermal resistance () is negligible, indicating perfect heat transfer across the interface. On the other hand, at high interfacial resistance ( K/W), heat transfer between the tip and sample becomes significantly suppressed, effectively isolating the thermal transport pathways. The resulting probe temperature profiles for different values are presented in Fig. 5(a). At high interface thermal resistance values (), the sample temperature remains constant at room temperature, indicating the absence of local heating. Therefore, cannot exceed , since local heating of the sample is need for thermal resistance measurement.
The effect of the on local sample heating was examined for two materials with strongly contrasting intrinsic thermal conductivities: Cu (300 W/mK) and SiO2 (1.4 W/mK). As shown in Fig. 4(b–d) for SiO2 and Fig. 4(f–h) for Cu, identical values of result in qualitatively similar surface temperature distributions beneath the tip, despite the large difference in bulk thermal conductivity. Because the COMSOL model considers smooth surfaces for all domains, the resulting temperature profiles for Cu and SiO2 are expected to be identical, which is in agreement with previous studies showing that is independent of the underlying material in the absence of surface roughness effects li2025quantitative; pernot2021frequency; bahrami2004thermal. At lower interfacial thermal resistance ( K/W), efficient heat transfer from the probe leads to a broader thermal diffusion profile in both materials. In contrast, increasing K/W strongly suppresses heat flow across the interface, resulting in negligible local surface heating for both Cu and SiO2. Importantly, the experimentally measured total thermal resistance lies in the range K/W and is accompanied by clear local surface heating beneath the probe. Therefore, must be K/W and cannot exceed the experimentally inferred thermal resistance.
Furthermore, the difference between the SThM tip temperature and the corresponding local sample temperature variation is plotted in Fig. 5(b). The change in tip and sample temperature () is shown for various values. The results shown in Fig. 5(a,b) demonstrate that plays a dominant role in the quantitative determination of the sample’s local thermal conductivity. While is treated parametrically in the FEM simulations to examine its influence on the measured temperature response, this approach is adopted to establish a physically constrained relationship between the experimentally accessible temperature contrast and . By comparing the experimentally measured (Fig. 3) with the trends predicted by the FEM model shown in Fig. 5 (b), a plausible and bounded range of is identified for the given probe–sample contact conditions. Although the thermal conductivity extraction model used in this study does not explicitly depend on the absolute value of , the FEM model serves as a sanity check and establishes a physically reasonable range for .
@@ IV.2 Experimental determination of model parameters
To experimentally validate the quantitative SThM model, a 15 nm aluminum (Al) thin film deposited on a SiO2 substrate was selected as the test sample. The film was patterned adjacent to the bare SiO2 substrate, enabling direct comparison between regions of differing thermal conductivity under identical measurement conditions. The sample preparation and fabrication details are provided in the Sec. III (A). All SThM measurements were carried out at a controlled room temperature of about 294 K, monitored continuously by a temperature regulation system. The SThM probe was positioned near the Al/SiO2 interface and brought into mechanical contact with the sample surface. This contact led to a sudden reduction in probe temperature, indicating the activation of an additional thermal conduction pathway at the tip–sample junction, thereby enhancing heat dissipation into the sample (see Fig. S1 from the Supplementary Material). Subsequently, as the tip scanned the sample area, the probe recorded thermal signals from the sample as temperature variations, which allowed the calculation of the sample’s thermal resistance. A total of 7 independent SThM scans were acquired to ensure the reproducibility of the measurements. The corresponding pixel-level thermal signal and topography data were extracted using the SPM image analysis software named Gwyddion. The data were processed and analyzed using a custom-developed Python script designed specifically for quantitative pixel-wise thermal signal correlation and noise reduction. This in-house code enabled automated data extraction, normalization, and statistical comparison across multiple scans, significantly improving the accuracy and consistency of the thermal mapping results.
@@
Figure. 6(a,c) shows the topography and thermal signal map of 15 nm-thick Al thin film patterned on a SiO2 substrate. A strong thermal signal contrast was observed between the Al film and the SiO2 substrate, reflecting the significant differences in their thermal conductivities. Two approach-retraction curves for either side of the interface are used for the conversion between the SThM signal and thermal resistance using Eq. (10). Further details about the approach-retraction curve are provided in Fig. S1 (Supplementary Material). The relative change in the thermal signal (20 mV) from the SiO2 surface to the Al film corresponds to a decrease in thermal resistance from the SiO2 surface to the 15 nm-thick Al film. Therefore, this decrease in thermal resistance indicates a higher thermal conductivity of the Al film compared to the SiO2 substrate.
@@
The thermal conductivity of Al is determined using the procedure described in Refs. huang2024violation and spiece2021quantifying, which reduces noise and measurement uncertainty. To extract the thermal resistance associated with Al and SiO2, a pixel-to-pixel correlation between the height and thermal signal maps is performed. This correlation is used to construct a 2D histogram (see Fig. S2 from the Supplementary Material), which enables a separation of the thermal responses from Al and SiO2 surfaces. Traditional analysis methods often rely on arbitrarily selected line-cuts or cropped regions of interest, introducing potential user bias in the selection process. In contrast, the present approach systematically processes all available data points from the thermal signals measurement, ensuring reproducibility and minimizing subjective interpretation. This comprehensive analysis leverages the full spatial information content, thereby improving statistical confidence in the extracted thermal properties. Subsequently, the acquired data are subjected to statistical processing to extract thermal properties. However, this analysis is susceptible to errors arising from experimental artifacts—such as surface contamination, tip-sample contact variations, or instrumental noise—which can distort the thermal signal and compromise measurement accuracy. Therefore, to quantitatively extract representative thermal resistance values while accounting for spatial heterogeneity, a 2D histogram of thermal resistance on the y-axis and film thickness on the x-axis is constructed (Fig. S2 of Supplementary Material). This histogram is then projected onto the thermal resistance axis to obtain statistically representative values. This projection combines measurements across all thickness values, yielding a one-dimensional distribution from which statistically robust resistance values can be determined. The resulting 1D distribution is modeled using a double-Gaussian function, from which the mean peak positions corresponding to the thermal resistances of Al and SiO2 are extracted. The statistical uncertainties in these mean values are obtained from the covariance matrix of the nonlinear least-squares fitting procedure. From this analysis, the mean thermal resistances were found to be 2.60 0.03 MK/W and 5.41 0.04 MK/W for Al and SiO2, respectively.
@@
In the following analysis, the thermal resistance distributions presented in Figure 6 (e,f) are utilized to quantitatively determine the thermal conductivity of 15 nm Al thin film. Following the theoretical framework established with Eq. (13) in Sec. III, we model the measured thermal resistance as a series combination of the tip thermal resistance (), sample spreading resistance (), and the film-substrate interfacial thermal resistance () as previously described in Ref. [huang2024violation]. This formulation enables direct extraction of the Al thermal conductivity from the experimental data. Therefore, the measured thermal resistance () consists of three components: , , and . Among these contributions, as calculated in Sec. III(C), K/W, which is at least two orders of magnitude smaller than both and . Furthermore, because the thermal conductivity is evaluated relative to the SiO2 reference surface, the contribution of can be neglected.
Furthermore, since the experimentally measured thermal resistance () values are K/W, it can be concluded that is also K/W, as discussed in Sec. IV(A). In addition, the tip–sample contact resistance () is mainly governed by surface roughness, the applied tip–sample normal force, and the resulting contact area pernot2021frequency; bahrami2004thermal; puyoo2010thermal. Therefore, the thermal contact between the SThM probe and the sample may slightly differ for Al and SiO2 surfaces; however, it is very challenging to determine the exact for each case. Therefore, for simplification, we assumed the same for both Al and SiO2 interfaces. This assumption introduces negligible uncertainty in the measurement because both surfaces exhibit comparable roughness (see Fig. S6 in the Supplementary Material). Additionally, the tip-sample normal contact force was maintained with an AFM setpoint value of 0.6 V, which corresponds to < 10 nN contact force for this type of probes. This assumption allows us to isolate the spreading resistance component, which leads to the measured thermal resistance () on the Al film surface. Finally, the thermal conductivity is calculated using Eq. (15), under the assumption of a semi-infinite substrate. The analysis of the statistical distribution of thermal resistance values yields W/mK for the 15 nm Al film. Detailed calculation of is provided in S1 of Supplementary Material. We assumed a literature value for the thermal conductivity of 1.38 W/mK for the SiO2 substrate grove1967physics. The reported uncertainty accounts for both statistical variations and thermal signal errors arising from additional heat transfer through the water meniscus formed around the SThM tip assy2015temperature.
V Discussion
@@
The thermal conductivity of Al consists of both phonon () and electron () contributions, with electrons typically dominating heat transport in bulk. Jain et al. performed detailed first-principles calculations to examine both phonon-phonon and electron-phonon scattering processes in bulk Al, quantitatively determining the individual contributions of and to the overall thermal conductivity jain2016thermal. This study further quantified the mean free path () distributions of both phonon and electron heat carriers by analyzing their respective thermal conductivity accumulation functions at 300 K (see Fig. S7 of Supplementary Material). Analysis of the thermal conductivity accumulation function demonstrates that the electron contribution decreases substantially as decreases, indicating that electrons with short contribute less to heat transport than those with longer . This suggests that long-range electron transport dominates thermal conductivity in bulk Al, whereas phonon transport shows the inverse behavior, being dominated by the short scattering processes.
@@
The accumulation function data at 300 K indicate that phonon thermal conductivity in Al is dominated by carriers with mean free paths ranging from 1 nm to 22 nm (comprising of ), while electron thermal conductivity requires longer mean free paths of 10 nm to 25 nm for similar accumulation ( of ). The onset of contributions also differs markedly: phonon transport begins at , whereas electron transport only becomes substantial at . These characteristic length scales directly impact thermal transport in confined geometries. In our 15 nm Al film, a sub-stoichiometric Al2O3-x layer (1–2 nm thick) forms naturally on the Al surface due to exposure to ambient air yang2018liquid; gorobez2021growth, reducing the effective conductive metallic thickness to approximately 13–14 nm. This confinement preferentially suppresses electron transport, as the film dimension is smaller than the typical electron mean free paths, resulting in diffuse boundary scattering that significantly reduces . Conversely, phonons with mean free paths of 1–13 nm can propagate through the film with less boundary interference. This size-dependent scattering leads to a crossover in the dominant heat carrier: while bulk Al thermal conductivity is electron-dominated, phonon transport becomes the primary mechanism in this nanoscale thin film. Therefore, the thermal conductivity obtained in this study ( W/mK) falls within a reasonable and expected range for nanoscale Al films. This value, approximately 80 times lower than bulk Al, is consistent with the dominance of phonon transport and the suppression of electron contributions due to dimensional confinement.
@@
The thermal signal and topography appear elevated along the Al/SiO2 boundary (see Fig. 6(a, c)). The elevated topography introduces a minor artifact in the thermal signal due to AFM feedback sensitivity effects at the step height. Consequently, the higher thermal resistance extracted at the edge arises from the same artifact, as it is directly derived from the thermal signal. Nevertheless, Fig. 6(e) demonstrates the significant potential of this approach for probing thermal transport variations across dislocation defects and grain boundaries with sub-100 nm spatial resolution. Furthemore, validation using established approaches, such as those described in Refs. tiwari2013anisotropic; isotta2023microscale for assessing thermal property variations across grain boundaries with different densities, would be highly beneficial, particularly given the high spatial resolution offered by SThM.
VI Conclusion